Problem 2 (1 point) A slot machine has 4 slots. After the push of a button, each slot on the machine will show one of the three following pictures: a cherry, a lemon, or a star. A: exactly 1 cherry(s) appears B: exactly 2 lemon(s) appears C: at least one non-fruit appears Part (a) How many elements are in this sample space? There is no need to list them. (If rounding use at least four digits after the decimal Part (b) Find P(B) (If rounding use at least four digits after the decimal in your answer)stat
Answer #1

a) Given a slot machine has 4 slots and after the push of a button each slot may show the three pictures of a cherry, a lemon or a star.

Therefore the number of elements in the sample space are 3*3*3*3=81

b) Let B be the event that exactly two lemons appears.

therefore n(B)=6

P(B)=6/81

=2/27